# SAINT: Attention-Based Modeling of Sub-Action Dependencies in Multi-Action Policies

### v1

@@ A.2 Learning Curves Including Wol-DDPG
Figure 5 presents the full training curves for all methods in the CityFlow environments, including Wol-DDPG. As noted in Section 5.1, Wol-DDPG consistently underperforms relative to other methods. This poor performance is consistent with prior findings [5], which show that Wol-DDPG is ill-suited to settings with unordered sub-actions. Learning curves excluding Wol-DDPG are presented in Figure 2.
@@ B.1 Environmental Setup
The Combinatorial Navigation Environment (CoNE) [20] is designed to evaluate RL algorithms in settings with high-dimensional, combinatorial action spaces and strong, state-dependent sub-action dependencies. In CoNE, actions are formed by simultaneously selecting discrete sub-actions, each specifying movement along a different dimension. These sub-actions are executed in parallel to produce a composite transition, which may advance the agent toward the goal or result in failure by entering a pit.
@@
To our knowledge, no other existing benchmarks offer the combination of large-scale action spaces and tunable sub-action dependencies found in CoNE. Popular environments such as the DeepMind Control Suite [37], for example, lack meaningful sub-action interactions [3].
@@ Appendix C State Conditioning
Appendix D SAINT’s Computation Cost

### v2

@@ A.2 Learning Curves Including Wol-DDPG
Figure 5 presents the full training curves for all methods in the CityFlow environments, including Wol-DDPG. As noted in Section 5.1, Wol-DDPG consistently underperforms relative to other methods. This poor performance is consistent with prior findings Chen et al. (2023), which show that Wol-DDPG is ill-suited to settings with unordered sub-actions. Learning curves excluding Wol-DDPG are presented in Figure 2.
@@ B.1 Environmental Setup
The Combinatorial Navigation Environment (CoNE) Landers et al. (2024) is designed to evaluate RL algorithms in settings with high-dimensional, combinatorial action spaces and strong, state-dependent sub-action dependencies. In CoNE, actions are formed by simultaneously selecting discrete sub-actions, each specifying movement along a different dimension. These sub-actions are executed in parallel to produce a composite transition, which may advance the agent toward the goal or result in failure by entering a pit.
@@
To our knowledge, no other existing benchmarks offer the combination of large-scale action spaces and tunable sub-action dependencies found in CoNE. Popular environments such as the DeepMind Control Suite Tassa et al. (2018), for example, lack meaningful sub-action interactions Beeson et al. (2024).
@@ Appendix C State Conditioning
Appendix D Comparison to CTDE Architecture
Centralized Training, Decentralized Execution (CTDE) is a widely used paradigm in cooperative multi-agent RL (Iqbal and Sha, 2019; Rashid et al., 2020; Yu et al., 2022). It is motivated by settings wherein execution is inherently decentralized — agents must act from local information (e.g., due to communication or latency constraints) — while training can exploit globally aggregated experience. CTDE is particularly well-suited to problems with heterogeneous observations and truly distinct decision-makers.
In the combinatorial action settings we consider, control is frequently centralized — a single policy outputs the entire joint action. In this regime, the “agents” are not distinct learners, so the usual CTDE motivation applies only weakly. Framing sub-actions as “agents” is therefore simply a modeling choice that imposes a factorized execution-time policy:
$\pi(\mathbf{a}\mid s)=\prod_{i}\pi_{i}(a_{i}\mid o_{i})\;,$
for some choice of per-sub-action observations $o_{i}$. Under this construction, CTDE largely evaluates whether providing global information to the value function can offset the restriction to a factorized policy at execution time, rather than directly addressing how to represent dependencies in the joint-action distribution.
For completeness — and because CityFlow admits a natural multi-agent interpretation with heterogeneous local observations across intersections, particularly in the Irregular setting — we report a CTDE baseline, MAPPO (Yu et al., 2022).
Figure 9: Comparison to MAPPO on CityFlow Irregular. MAPPO underperforms the fully factorized policy and exhibits greater training instability, indicating that centralized critic information is insufficient to induce coordinated behavior. Curves are averaged over 5 seeds; shaded regions show $\pm 1$ standard deviation.
As shown in Figure 9, this CTDE baseline does not outperform the fully factorized policy and exhibits greater training instability. These results indicate that, even in a setting wherein CTDE is most well-motivated, centralized critic information alone is insufficient; policy representations must also capture joint-action dependencies.
Appendix E Robustness to Offline RL Training Objective
The results in Section 5.4 show that SAINT’s set-based, permutation-invariant architecture learns effective representations for offline RL in combinatorial action spaces when trained with an BCQ objective. To assess whether these advantages persist across different offline RL methods, we also evaluate SAINT with two additional objectives, Advantage Weighted Actor Critic (AWAC) (Nair et al., 2020) and Implicit Q-learning (IQL) (Kostrikov et al., 2021). These experiments use the same medium-expert datasets considered in Section 5.4 and follow an identical controlled protocol: for each algorithm, we instantiate factorized, autoregressive, and SAINT-based policy parameterizations while keeping the critic, training procedure, and hyperparameters fixed.
E.1 AWAC
Task SAINT Factored AR
[…]
Average 676.5 $511.0$ $520.5$
Table 5: Mean $\pm$ std performance on offline DM Control tasks with AWAC variants.
E.2 IQL
Task SAINT Factored AR
[…]
Average Return 709.5 $676.2$ $684.9$
Table 6: Mean $\pm$ std performance on offline DM Control tasks with IQL variants.
Tables 5 and 6 show that SAINT achieves the strongest performance across all domains under both AWAC and IQL, consistent with the BCQ results in the main text. This indicates that the benefits of modeling sub-action interactions extend across offline RL objectives and are not tied to a specific learning algorithm.
Appendix F SAINT’s Computation Cost
@@
Appendix G Robustness to Architectural Hyperparameters
We systematically evaluated SAINT’s sensitivity to architectural hyperparameters by sweeping over the number of self-attention blocks $\{1,3,5\}$ and attention heads $\{1,2,4,8\}$, yielding twelve configurations.
Configuration Mean Return
1 block $\times$ 1 head $13376.6\pm 973.3$
1 block $\times$ 2 heads $13416.6\pm 1003.6$
1 block $\times$ 4 heads $13058.6\pm 725.5$
1 block $\times$ 8 heads $\mathit{-13744.5\pm 900.6}$
3 blocks $\times$ 1 head $\mathbf{-12834.6\pm 499.8}$
3 blocks $\times$ 2 heads $13195.0\pm 618.8$
3 blocks $\times$ 4 heads $13209.5\pm 778.8$
3 blocks $\times$ 8 heads $13156.5\pm 581.9$
5 blocks $\times$ 1 head $13320.7\pm 831.7$
5 blocks $\times$ 2 heads $13658.8\pm 855.6$
5 blocks $\times$ 4 heads $13664.3\pm 961.5$
5 blocks $\times$ 8 heads $13664.8\pm 822.3$
Factored PPO $14200.5\pm 1127.4$
AR PPO $13995.9\pm 789.4$
Standard PPO $14442.5\pm 1180.0$
Table 7: Mean episodic return $\pm$ standard deviation on CityFlow Irregular. We varied the number of attention blocks $\{1,3,5\}$ and the number of attention heads $\{1,2,4,8\}$, for a total of 12 configurations. All SAINT variants outperform Factored, AR, and PPO baselines. The best SAINT configuration is in bold, the worst is in italics.
A consistent pattern emerges in Table 7. Moderate depth (3 blocks) with 2–4 heads yields strong and stable performance, while very high head counts (8) tend to degrade results. Crucially, every SAINT variant outperforms all baselines, including Factored PPO, AR-PPO, and standard PPO. This robustness implies that SAINT’s architectural advantages are not narrowly tied to a specific hyperparameter regime but instead generalize across a broad design space. Careful tuning can yield an additional 5–10% improvement, yet even suboptimal choices consistently achieve better outcomes than state of the art methods.